x              y                  xy             x2                      y2

3              4                 12              9                  16

4              6                 24             16                 36

5              7                35               25                49

7             12                84              49               144

8            14                112             64                196

_____________________________________________

27           43               267             163             441

Find the slope and intercept
Answer #1
X Y X * Y X2
3 4 12 9
4 6 24 16
5 7 35 25
7 12 84 49
8 14 112 64
Total 27 43 267 163

Equation of regression line is Ŷ = a + bX
b = ( n Σ(XY) - (ΣX* ΣY) ) / ( n Σ X2 - (ΣX)2 )
b = ( 5 * 267 - 27 * 43 ) / ( 5 * 163 - ( 27 )2)
b = 2.0233

a =( ΣY - ( b * ΣX ) ) / n
a =( 43 - ( 2.0233 * 27 ) ) / 5
a = -2.3256
Equation of regression line becomes Ŷ = -2.3256 + 2.0233 X

Slope  b = 2.0233

Intercept a = - 2.3256